# A Marketplace for AI-Generated Adult Content and Deepfakes

### v1

@@ 1 Introduction
Text-to-image generative AI has fundamentally transformed image creation, enabling anyone with internet access to produce highly realistic images from simple text descriptions. Modern diffusion models such as Stable Diffusion Podell et al. (2023), DALL-E Ramesh et al. (2021), and Midjourney allow users to create everything from artistic illustrations and product mockups to photorealistic portraits and scenes that are increasingly difficult to distinguish from authentic photographs Corvi et al. (2023); Cazenavette et al. (2024).
@@
The rapid adoption of these models has democratized visual content generation, giving rise to a new class of online platforms that blend social interaction with AI-powered content creation. Platforms such as Civitai, [^1] https://civitai.com PixAI, [^2] https://pixai.art/en and TensorArt [^3] https://tensor.art/ have become popular hubs where users gather around shared tools, models, and generated outputs. Users in these new environments simultaneously act not only as creators and consumers of AI-generated media, but also of the underlying generative tools themselves. These platforms therefore constitute a distinct digital space, raising new questions about their governance and content moderation, as well as the social implications of widespread access to powerful generative technologies Thiel et al. (2024).
@@
A growing body of research highlights how these platforms can also facilitate the production of harmful or misleading content. Generative AI has been used to create political disinformation, identity fraud, child sexual abuse material, and non-consensual intimate imagery, often at low cost and high scale Fu et al. (2023); Wei et al. (2024); Yang et al. (2024); Hawkins et al. (2025); Ciardha et al. (2025). The ease with which such content can be generated and circulated exacerbates longstanding challenges around consent, authenticity, and accountability, motivating closer empirical scrutiny of how AI-centric platforms structure participation and enforce safeguards Thiel et al. (2024).
@@
Against this broader landscape, the Civitai platform provides a particularly useful case for studying governance and incentives within these social generative AI ecosystems. In the United States, Civitai has rapidly grown into the largest AI-centric content platform, reaching approximately 3 million registered users and 12–13 million unique monthly visitors by November 2024. Founded in 2022, the platform has raised over US\$5 million in venture funding, including investment from Andreessen Horowitz, signaling increasing mainstream interest in these platforms Perez (2023). At the same time, Civitai’s open structure has drawn criticism for enabling intellectual property violations Yamamoto et al. (2024) and facilitating the circulation of child sexual abuse material and non-consensual deepfakes Thiel et al. (2024, 2023); Epstein et al. (2023); Maiberg (2023a); Hawkins et al. (2025). These concerns echo earlier warnings about the societal effects of generative media Karnouskos (2020); Al-Khazraji et al. (2023); Hancock and Bailenson (2021) and align with mounting evidence of increasing online harassment and gendered harm linked to deepfake technologies Amin (2025); Lazard et al. (2025); Akter and Ahmed (2025).
@@
Civitai differs from other generative AI platforms not only in scale, but also in platform design. While platforms such as PixAI and TensorArt primarily center on sharing, remixing, and showcasing AI-generated content, Civitai uniquely incorporates a formal bounty system that allows users to post paid requests for specific types of AI-generated media Maiberg (2023b). Figure 1 presents two example bounties, including their titles, descriptions, and reference images. Each bounty is posted by a user (referred to as a Supporter) who offers virtual currency (buzz) to reward one or more winning submissions. Buzz can be purchased (1,000 buzz per US dollar) and spent within the platform. Bounties progress through several states—Active, Awarded, or Expired—and user engagement is reflected through the number of likes, comments, and submitted entries. Through this system, users who lack the time or technical expertise to generate desired content can offer monetary rewards for specific requests, creating incentives for other users to compete to produce and submit entries. This feature introduces a structured marketplace for AI-generated content that directly coordinates demand, labor, and reward—an incentive mechanism that is absent from comparable platforms.
@@
Existing empirical studies of Civitai have begun to characterize the content produced on the platform. (Wei et al., 2024) analyze the themes, prompts, and engagement patterns of abusive images and models, finding that creators of such content often gain increased prominence. (Palmini et al., 2024) introduce a massive dataset of Civitai prompts, revealing a strong preference for explicit content and raising concerns about the reproduction of harmful stereotypes. Yet these studies focus primarily on generated images and models themselves, leaving the dynamics of Civitai’s bounty marketplace largely unexplored.
@@ 4.3 Gendered targeting and ethical concerns around deepfakes
These patterns align with broader evidence that deepfake technologies disproportionately harm women and are frequently used in non-consensual or sexually exploitative contexts Laffier and Rehman (2023); Kira (2024); Lazard et al. (2025). Even when requests are classified as SFW, they may still involve the generation of images without the knowledge or consent of the targeted individual.

### v2

@@ 1 Introduction
Text-to-image generative AI has fundamentally transformed image creation, enabling anyone with internet access to produce highly realistic images from simple text descriptions. Modern diffusion models such as Stable Diffusion Podell et al. [2023], DALL-E Ramesh et al. [2021], and Midjourney allow users to create everything from artistic illustrations and product mockups to photorealistic portraits and scenes that are increasingly difficult to distinguish from authentic photographs Corvi et al. [2023]; Cazenavette et al. [2024].
@@
The rapid adoption of these models has democratized visual content generation, giving rise to a new class of online platforms that blend social interaction with AI-powered content creation. Platforms such as Civitai, [^1] https://civitai.com PixAI, [^2] https://pixai.art/en and TensorArt [^3] https://tensor.art/ have become popular hubs where users gather around shared tools, models, and generated outputs. Users in these new environments simultaneously act not only as creators and consumers of AI-generated media, but also of the underlying generative tools themselves. These platforms therefore constitute a distinct digital space, raising new questions about their governance and content moderation, as well as the social implications of widespread access to powerful generative technologies Thiel et al. [2024].
@@
A growing body of research highlights how these platforms can also facilitate the production of harmful or misleading content. Generative AI has been used to create political disinformation, identity fraud, child sexual abuse material, and non-consensual intimate imagery, often at low cost and high scale Fu et al. [2023]; Wei et al. [2024]; Yang et al. [2024]; Hawkins et al. [2025]; Ciardha et al. [2025]. The ease with which such content can be generated and circulated exacerbates longstanding challenges around consent, authenticity, and accountability, motivating closer empirical scrutiny of how AI-centric platforms structure participation and enforce safeguards Thiel et al. [2024].
@@
Against this broader landscape, the Civitai platform provides a particularly useful case for studying governance and incentives within these social generative AI ecosystems. In the United States, Civitai has rapidly grown into the largest AI-centric content platform, reaching approximately 3 million registered users and 12–13 million unique monthly visitors by November 2024. Founded in 2022, the platform has raised over US\$5 million in venture funding, including investment from Andreessen Horowitz, signaling increasing mainstream interest in these platforms Perez [2023]. At the same time, Civitai’s open structure has drawn criticism for enabling intellectual property violations Yamamoto et al. [2024] and facilitating the circulation of child sexual abuse material and non-consensual deepfakes Thiel et al. [2024, 2023]; Epstein et al. [2023]; Maiberg [2023a]; Hawkins et al. [2025]. These concerns echo earlier warnings about the societal effects of generative media Karnouskos [2020]; Al-Khazraji et al. [2023]; Hancock and Bailenson [2021] and align with mounting evidence of increasing online harassment and gendered harm linked to deepfake technologies Amin [2025]; Lazard et al. [2025]; Akter and Ahmed [2025].
@@
Civitai differs from other generative AI platforms not only in scale, but also in platform design. While platforms such as PixAI and TensorArt primarily center on sharing, remixing, and showcasing AI-generated content, Civitai uniquely incorporates a formal bounty system that allows users to post paid requests for specific types of AI-generated media Maiberg [2023b]. Figure 1 presents two example bounties, including their titles, descriptions, and reference images. Each bounty is posted by a user (referred to as a Supporter) who offers virtual currency (buzz) to reward one or more winning submissions. Buzz can be purchased (1,000 buzz per US dollar) and spent within the platform. Bounties progress through several states—Active, Awarded, or Expired—and user engagement is reflected through the number of likes, comments, and submitted entries. Through this system, users who lack the time or technical expertise to generate desired content can offer monetary rewards for specific requests, creating incentives for other users to compete to produce and submit entries. This feature introduces a structured marketplace for AI-generated content that directly coordinates demand, labor, and reward—an incentive mechanism that is absent from comparable platforms.
@@
Existing empirical studies of Civitai have begun to characterize the content produced on the platform. [Wei et al., 2024] analyze the themes, prompts, and engagement patterns of abusive images and models, finding that creators of such content often gain increased prominence. [Palmini et al., 2024] introduce a massive dataset of Civitai prompts, revealing a strong preference for explicit content and raising concerns about the reproduction of harmful stereotypes. Yet these studies focus primarily on generated images and models themselves, leaving the dynamics of Civitai’s bounty marketplace largely unexplored.
@@ 4.3 Gendered targeting and ethical concerns around deepfakes
These patterns align with broader evidence that deepfake technologies disproportionately harm women and are frequently used in non-consensual or sexually exploitative contexts Laffier and Rehman [2023]; Kira [2024]; Lazard et al. [2025]. Even when requests are classified as SFW, they may still involve the generation of images without the knowledge or consent of the targeted individual.